OpenStudy (genny7):

Please help! 4/7 + 1/8 + 1/2

OpenStudy (anonymous):

I got 1.2

OpenStudy (genny7):

Well i got 1 5/28 but i checked to see if thats the answer and its not

OpenStudy (anonymous):

you can do it directly by calculator or you simplify it first by choosing a CF fir the three fractions and i see 56 is a good CF

OpenStudy (genny7):

I actually simplified & i got 56 for the CF @esamalaa

OpenStudy (anonymous):

good what is your result?

OpenStudy (kailee1423):

I did it by calculator and got 1 and 11/56 with 1 as the whole number

OpenStudy (genny7):

Okay so this is what i did after i simplified i mutliplied them & i got 32/56 + 4/56 + 28/56 = 64/56 then i reduced it to 1 10/56 @esamalaa

OpenStudy (anonymous):

4/56? how you got that it should be 7/56

OpenStudy (genny7):

Oh now i see my mistake the 4 shouldnt be there...instead the 7 should lol @esamalaa

OpenStudy (anonymous):

yes

OpenStudy (genny7):

Can you check my work for another problem? @esamalaa

OpenStudy (genny7):

Problem: 3/5 + 1/8 + 1/8 My work: 24/45 + 5/45 + 5/45 = 34/45

OpenStudy (anonymous):

why you choose 45 ? and even you chosen 45 these results are not right

OpenStudy (genny7):

I was looking at it and its supposed to be 40, right?

OpenStudy (genny7):

@esamalaa

OpenStudy (anonymous):

yes right

OpenStudy (genny7):

Okay i think I got it from there...thanks for your help :) @esamalaa

OpenStudy (anonymous):

anytime ^_^

OpenStudy (genny7):

^_^